OpenStudy (anonymous):

Write the equation of the line that passes through (-1, 4) with a slope of -1 in standard form

OpenStudy (linknissan):

(y-4)=-1(x-(-1)) y-4=-1(x+1) y-4=-x-1 x+y=3

OpenStudy (anonymous):

thanks!(:

OpenStudy (linknissan):

anytime..nicolee!!